OpenStudy (anonymous):

Find the value of x. logx 1/1000000=−6

OpenStudy (anonymous):

\[\log_{x} 1/1000000=-6 \]

OpenStudy (anonymous):

log3 9 = 2 means 3^2 = 9 logx 9 = 2 means x^2 = 9 x = sqrt (9) x = 3, try it :D

OpenStudy (anonymous):

didnt work tried it

OpenStudy (anonymous):

the answer is 10

OpenStudy (anonymous):

10 ^ -6 = 1/10^6 = 1/1000000

OpenStudy (anonymous):

i got it.. thanks...

OpenStudy (anonymous):

medal for me :D

OpenStudy (anonymous):

done